# Interpret the estimand framework from a causal inference perspective

### v1

@@ 3.4 The while on treatment strategy
The while on treatment strategy is to use available endpoints measured before intercurrent events and discard any data afterwards. In case (3), participants survive the first assessment but the endpoint is not measured at the second assessment due to death. Through the while on treatment strategy, we use the first assessment on the endpoint and all data up to the first assessment when participants from case (3) are included in primary analysis. This means that, for participants with intercurrent events, we only use their data that are recorded during treatment before intercurrent events. The estimand formula is still $E(Y_{i}(X_{i}(R_{i}=1)=1)-Y_{i}(X_{i}(R_{i}=0)=0)~{}|~{}C)$, but now the definition of $Y$ has been changed from measurement at the second assessment to measurement at the first or second assessment.
@@ 3.5 The principal stratum strategy
In case (4), suppose we only want to know the treatment effect in the participants who would tolerate treatment toxicity. Tolerability of treatment toxicity is an intrinsic characteristic and should be interpreted from a hypothetical perspective. Suppose tolerability is represented by $T$. $T=0$ represents not tolerating treatment and $T=1$ represents tolerating treatment. Now, there are four subpopulations represented by $P$. $P=1$ are participants who would tolerate both the control and experimental treatments if they took both treatments. $P=2$ are participants who would tolerate neither the control treatment nor the experimental treatment if they took both treatments. $P=3$ are participants who would tolerate the control treatment and would not tolerate the experimental treatment if they took both treatments. $P=4$ are participants who would not tolerate the control treatment and would tolerate the experimental treatment if they took both treatments. And the subpopulation of participants who would tolerate both treatments, that is $P=1$, becomes our new target population. These four subpopulations may not be directly observed in real world and may have to be estimated from real data. They are different from subgroups of participants who tolerate or do not tolerate treatment in real world. For example, there is a subgroup of participants in the control arm who tolerate the control treatment. From this subgroup, some participants might tolerate the experimental treatment if assigned to the treatment arm instead, while the others might not tolerate the experimental treatment if assigned to the treatment arm instead. Hence, this subgroup can be a mix of $P=1$ and $P=3$.
Potential outcomes should be modified to take $T$ into account. For example, $Y_{i}(X_{i}(R_{i}=0)=0,T_{i}=0)$ would be the outcome if the participant took but did not tolerate the control treatment. $Y_{i}(X_{i}(R_{i}=1)=1,T_{i}=0)$ would be the outcome if the participant took but did not tolerate the experimental treatment. $Y_{i}(X_{i}(R_{i}=0)=0,T_{i}=1)$ would be the outcome if the participant took and tolerated the control treatment. $Y_{i}(X_{i}(R_{i}=1)=1,T_{i}=1)$ would be the outcome if the participant took and tolerated the experimental treatment. The estimand formula now becomes $E(Y_{i}(X_{i}(R_{i}=1)=1)-Y_{i}(X_{i}(R_{i}=0)=0)~{}|~{}C,P=1)$, with no changes in the definitions of $X$ and $Y$.
This strategy estimates the pure treatment effect in a different target population. When specific subpopulations are of clinical interest, this strategy would be appropriate. Statistical methods with regard to this strategy may be complex, with more assumptions. Bayesian inference would be a good choice for this strategy.
@@ 3.6 Further possibilities
The above five strategies proposed in the estimand framework can be used in a flexible way. Several strategies can be used together. For example, the treatment policy strategy to analyze concomitant therapies is used with the principal stratum strategy to identify a subpopulation. A composite strategy that the treatment policy strategy to analyze allowed concomitant therapies is combined with the hypothetical strategy to analyze prohibited concomitant therapies has been proposed by industrial professionals [4].
In addition to the five strategies, there are also other strategies to analyze intercurrent events. For example, in case (1), suppose the dose of the control or experimental treatment has to be reduced after concomitant therapies are used and concomitant therapies affect the endpoint. We define a new variable $M$ to represent concomitant therapies. $M$ can be a binary variable that indicates whether concomitant therapies are used or not. It can also be the total dose or the dose frequency of concomitant therapies. Hence, $M$ is a new confounder that affects both $X$ and $Y$. And figure 3 is a causal directed acyclic graph that shows the causal relationships between $X$, $Y$, $R$, $C$, $M$.
@@
The estimand formula now should be conditional on $M$ as $E(Y_{i}(X_{i}(R_{i}=1)=1)-Y_{i}(X_{i}(R_{i}=0)=0)~{}|~{}C,M)$, but there is no change in the meaning of the estimand. Compared to the treatment policy strategy and the hypothetical strategy, this strategy of defining intercurrent events as confounders and using model adjustment does not modify the definitions of $X$, $Y$ and the ATE, or modify any trial data. This strategy would be appropriate if the confounding effects of intercurrent events can be well adjusted for.
@@
4 Target populations versus analysis sets
Analysis sets are a group of participants that meets specific criteria and is used for specific analytical objectives. ICH E9 proposed different analysis sets, including full analysis set (FAS) and per protocol set (PPS) [2]. There are also other analysis sets, including intention to treat set (ITTS) and safety set (SS). ITTS is usually all participants enrolled in study. SS comes from ITTS and is usually a group of participants who take treatment at least once. FAS comes from SS and excludes some participants whose data are not suitable for analysis, such as those who take no treatment after randomization. PPS comes from FAS and only includes participants whose data have no or little protocol deviation.
Different analysis sets can lead to different treatment effect estimates, because from the estimand framework, they are different target populations. In the estimand framework, analysis sets are actually not required. We can define analysis sets into target populations or use the principal stratum strategy to estimate treatment effects in subpopulations equal to analysis sets. However, analysis sets are usually subsets of all trial data. If we build statistical models through the estimand framework on data subsets, we may introduce selection bias and break randomization of the assignment scheme, which means that the participants in an analysis set may not be considered randomized any more. Without the randomization assumption, statistical inference on estimands may be unviable or biased. Hence, we should carefully examine randomization in analysis sets if they are used as target populations.

### v2

@@ 3.4 The while on treatment strategy
The while on treatment strategy is to use available endpoints measured before intercurrent events and discard any clinical trial data afterwards. In case (3), participants survive the first assessment, but the endpoint is not measured at the second assessment due to death. Through the while on treatment strategy, we use the first assessment on the endpoint and all data up to the first assessment when participants are included in primary analysis. This means that, for participants with intercurrent events, we only use their data that are recorded during treatment before intercurrent events. The estimand formula is still E(Y(X(R=1)=1)Y(X(R=0)=0)|C)E(Y(X(R=1)=1)-Y(X(R=0)=0)~|~C), but now the definition of $Y$ has been changed from measurement at the second assessment to measurement at the first or second assessment.
@@ 3.5 The principal stratum strategy
In case (4), suppose we only want to know the treatment effect in the participants who would tolerate treatment toxicity. Tolerability of treatment toxicity is an intrinsic characteristic and should be interpreted from a hypothetical perspective. Suppose tolerability is represented by $T$. $T=0$ represents not tolerating treatment toxicity, and $T=1$ represents tolerating treatment toxicity. Now, there are four subpopulations represented by $P$. $P=1$ are participants who would tolerate both the control and experimental treatments if they took both treatments. $P=2$ are participants who would tolerate neither the control treatment nor the experimental treatment if they took both treatments. $P=3$ are participants who would tolerate the control treatment and would not tolerate the experimental treatment if they took both treatments. $P=4$ are participants who would not tolerate the control treatment and would tolerate the experimental treatment if they took both treatments. And the subpopulation of participants who would tolerate both treatments, that is $P=1$, becomes our new target population. These four subpopulations may not be directly observed in the real world and may have to be estimated from data. They are different from subgroups of participants who tolerate or do not tolerate treatment toxicity in the real world. For example, there is a subgroup of participants in the control arm who tolerate the control treatment. From this subgroup, some participants might tolerate the experimental treatment if assigned to the treatment arm instead, while the others might not tolerate the experimental treatment if assigned to the treatment arm instead. Hence, this subgroup may be a mix of $P=1$ and $P=3$.
Potential outcomes should be modified to take $T$ into account. For example, $Y_{i}(X_{i}(R_{i}=0)=0,T_{i}=0)$ would be the outcome if the participant took but did not tolerate the control treatment. $Y_{i}(X_{i}(R_{i}=1)=1,T_{i}=0)$ would be the outcome if the participant took but did not tolerate the experimental treatment. $Y_{i}(X_{i}(R_{i}=0)=0,T_{i}=1)$ would be the outcome if the participant took and tolerated the control treatment. $Y_{i}(X_{i}(R_{i}=1)=1,T_{i}=1)$ would be the outcome if the participant took and tolerated the experimental treatment. The estimand formula now becomes E(Y(X(R=1)=1)Y(X(R=0)=0)|C,P=1)E(Y(X(R=1)=1)-Y(X(R=0)=0)~|~C,P=1), with no changes in the definitions of $X$ and $Y$.
This strategy estimates the pure treatment effect in a different target population. When specific subpopulations are of clinical interest, this strategy would be appropriate. Statistical methods with regard to this strategy may be complex. Bayesian inference would be a good choice for this strategy [13, 14].
@@ 3.6 Further possibilities
The above five strategies proposed in the estimand framework can be used in a flexible way. Several strategies can be used together. For example, the treatment policy strategy to analyze concomitant therapies can be used with the principal stratum strategy to identify a subpopulation. A composite strategy has been proposed by industrial professionals that the treatment policy strategy to analyze allowed concomitant therapies is combined with the hypothetical strategy to analyze prohibited concomitant therapies [9].
In addition to the five strategies, there are also possible other strategies to analyze intercurrent events. For example, in case (1), suppose the dose of the control or experimental treatment has to be reduced after concomitant therapies are used, and concomitant therapies affect the endpoint. We define a new variable $M$ to represent concomitant therapies. $M$ may be a binary variable that indicates whether concomitant therapies are used or not. It can also be the total dose or the dose frequency of concomitant therapies. Now, $M$ becomes a new confounder that affects both $X$ and $Y$. Figure 3 is a causal directed acyclic graph that shows the causal relationships between $X$, $Y$, $R$, $C$, $M$.
@@
The estimand formula now should be conditional on $M$ as E(Y(X(R=1)=1)Y(X(R=0)=0)|C,M)E(Y(X(R=1)=1)-Y(X(R=0)=0)~|~C,M). Hence, the causal linear model would become $Y^{o}=\beta_{0}+\beta_{1}X^{o}+\beta_{2}C+\beta_{3}M+\varepsilon$, where $\beta_{1}$ equals to the ATE after adjusting for new confounding effects.
This new strategy defines intercurrent events as confounders and uses model adjustment. Compared to the treatment policy strategy and the hypothetical strategy, the model adjustment strategy does not modify the definitions of $X$, $Y$ and the target population, and it does not modify any clinical trial data. If the confounding effects from intercurrent events can be well adjusted for, the model adjustment strategy would be appropriate.
@@
4 Comparison between target populations and analysis sets
An analysis set is a group of participants that meet specific criteria and is used for specific analytical objectives. ICH E9 proposed different analysis sets, including full analysis set (FAS) and per protocol set (PPS) [1]. There are also other analysis sets, including intention to treat set (ITTS) and safety set (SS). ITTS is usually all participants enrolled in study. SS comes from ITTS and is usually a group of participants who take treatments at least once. FAS comes from SS and may exclude some participants whose data are not suitable for analysis. PPS comes from FAS and may only include participants whose data have no or little protocol deviation. In practice, definitions of different analysis sets may vary by clinical trial. Some statistical analysis plans may give special considerations to analysis sets based on study design. We may have to interpret analysis sets in the clinical trial context.
Analysis sets are usually subsets of all trial data. Strictly speaking, they are not necessarily the target population of interest. If we build statistical models on data subsets through the estimand framework, we may introduce selection bias and break randomization of the assignment scheme, which means that the participants in an analysis set may not be considered randomized any more. Without the randomization assumption, statistical inference on estimands may be unviable or biased. Hence, we should carefully examine randomization in analysis sets if we want to use them as target populations. On the other hand, the principal stratum strategy provides an approach to estimate treatment effects in various analysis sets, if it is feasible to define analysis sets into subpopulations.